Find dimensions and basis of solution space W for a homogeneous system.
x^2+2y+2z-s+3t=0
x+2y+3z+s+t=0
3x+6y+8z+s+5t=0
1
Expert's answer
2020-04-13T10:16:38-0400

The augmented matrix is


"\begin{bmatrix}
 1 & 2& 2 & -1& 3 \\ 1 & 2& 3 & 1& 1\\ 3 & 6& 8 & 1& 5
 
\end{bmatrix}"

Reduce the system to echelon form:


"R_3\to R_3-3R_1\implies
\begin{bmatrix}
 1 & 2& 2 & -1& 3 \\ 1 & 2& 3 & 1& 1\\ 0 & 0& 2 & 4& -4
 \end{bmatrix}"

"R_2\to R_2-R_1\implies
\begin{bmatrix}
 1 & 2& 2 & -1& 3 \\ 0 & 0& 1 & 2& -2\\ 0 & 0& 2 & 4& -4
 \end{bmatrix}"

"R_3\to R_3-2R_2\implies
\begin{bmatrix}
 1 & 2& 2 & -1& 3 \\ 0 & 0& 1 & 2& -2\\ 0 & 0& 0 & 0& 0
 \end{bmatrix}"

This corresponds to the system:


"z+2s-2t=0,x+2y+2z-s+3t=0"

The system in echelon form has 2 nonzero equations in five unknowns. Hence the system has "5-2=3" free variables which are "y,s,t". Thus,


"\dim W=3"

We obtain the basis for W:

i. "y=1,s=0,t=0"


"v_1=(-2,1,0,0,0)"

ii. "y=0,s=1,t=0"


"v_2=(5,0,-2,1,0)"

iii."y=0,s=0,t=1"


"v_3=(-7,0,2,0,1)"

The set "\{v_1,v_2,v_3\}" is a basis of the solution space W.